1. Savannah can paint a room in seven hours. Jordan can paint the same room in nine hours. How long does it take for both Savannah and Jordan to paint the room if they are working together?
2. It takes Shelby 2 hours and 3 minutes to cut the grass. Abigail needs 3 hours and 46 minutes to cut the grass. If they work together, how long will it take them to cut the grass?
thank you for advance

Answer by stanbon(75887)

2. It takes Shelby 2 hours and 3 minutes to cut the grass. Abigail needs 3 hours and 46 minutes to cut the grass. If they work together, how long will it take them to cut the grass?
--------
I'll do the messier one; follow the same pattern to work #1.
-----------------------
#1
Shelby Time = 123 min/job ; rate = 1/123 job/min
---------
Abigail time= 226 min/job ; rate = 1/226 job/min
--------------
Together time = x min/job ; rate = 1/x job/min
---------------
EQUATION:
rate + rate = rate together
1/123 + 1/226 = 1/x
349/(123*226) = 1/x
x = (123*226)/349 =79.65 minutes
x = 1 hr 19 min 39 seconds (Time to do the job together)
----------------
Cheers,
Stan H.